OpenStudy (anonymous):

a tug boat goes 160 miles upstream in 20 hours. return trip takes 5 hours. find the speed of the tu boat without a current, and the speed of the current

hero (hero):

@mason2

OpenStudy (anonymous):

160/5=32 I got?

OpenStudy (anonymous):

I got speed of boat=20, speed, and current 12 ???

hero (hero):

Using the formula d = rt, in this case r = b ± c b = speed of boat c = speed of current Upstream 160 = 20(b - c) 160/20 = b - c 8 = b - c Downstream 160 = 5(b + c) 160/5 = b + c 32 = b + c So we have the two equations: 8 = b - c 32 = b + c Add both equations together to get: 40 = 2b 40/2 = b 20 = b The speed of the boat is 20mph c = b - 8 c = 20 - 8 c = 12 The speed of the current is 12 mph

hero (hero):

Well, at least you understand the methods

OpenStudy (anonymous):

thank you very much! I did a similar problem, got it right, but I forgot how I did it so thank you. How do I give you a medal? I am new here

hero (hero):

Click "Good Answer"

OpenStudy (anonymous):

you mean "best response"?